OpenStudy (anonymous):

Which quadratic has a graph with x-intercepts of 6 and -6?

OpenStudy (precal):



OpenStudy (anonymous):

y = x2 - 6 B) y = x2 - 36 C) y = x2 + 36 D) y = x2 -12x + 36

OpenStudy (anonymous):

those are the options.

OpenStudy (precal):

y=x^2-36

OpenStudy (precal):

because y=(x+6)(x-6) x=-6 and x=6

OpenStudy (anonymous):

You have to find the zeros of each equation. And in this case, the zeros need to be -6 and 6. In order to find the zeros, set each equation equal to zero and solve for x

OpenStudy (anonymous):

thank you you guys.

OpenStudy (precal):

I left the correct graph posted for you

OpenStudy (precal):

anytime :)